OpenStudy (dls):

A person takes 2 seconds to reach his maximum speed of 18km/hr.What is the magnitude of average acceleration?

mathslover (mathslover):

oh k wait please

mathslover (mathslover):

a = (v-u)/t = 5 / 2 ms^{-2}

mathslover (mathslover):

\[\large{\frac{v-u}{t}=a}\] \[\large{\frac{5}{2}ms^{-2} =a}\]

OpenStudy (dls):

what is u?

mathslover (mathslover):

Initial speed = u

mathslover (mathslover):

v= final speed

OpenStudy (dls):

i mean the value of u!!

OpenStudy (dls):

how did you get v-u as 5?

mathslover (mathslover):

since u = 0 m/s and v = 18 km/hr = 18 * 5/18 m/s = 5 m/s

OpenStudy (dls):

why is u 0?

OpenStudy (dls):

nowhere mentioned,you cant do that

mathslover (mathslover):

hmn.. wait then

OpenStudy (dls):

he was already at some initial speed and then then finally he speeded up to 5m/s

mathslover (mathslover):

options given?

OpenStudy (dls):

answer is 2.5m/s^2

mathslover (mathslover):

check the question again @DLS

OpenStudy (dls):

its correct but..

OpenStudy (dls):

i did

mathslover (mathslover):

the process I did is wrong since u is not equal to 0

OpenStudy (dls):

yes

OpenStudy (xishem):

Average acceleration is defined as:\[a_{avg}=\frac{\Delta v}{\Delta t}=\frac{v_f-v_i}{t_f-t_i}\]We have no choice to assume that the initial velocity is 0, and the rest of the information is given in the problem, so\[a_{avg}=\frac{18\frac{km}{hr}-0\frac{km}{hr}}{2s-0s}=\frac{18 \frac{km}{hr}}{2s} \times \frac{1hr}{3600s} \times \frac{1000m}{1km}=2.5\frac{m}{s^2}\]

mathslover (mathslover):

@Xishem also took u = 0 ..

mathslover (mathslover):

that's it..

mathslover (mathslover):

NO CHOICE ... other than to assume u = 0 m/s

OpenStudy (xishem):

You have to assume that the initial velocity is 0 since there is no initial velocity given. Average acceleration is dependent on knowing initial velocity beforehand.

OpenStudy (dls):

just how can we assume lol

OpenStudy (dls):

its just not possible

OpenStudy (dls):

logically

mathslover (mathslover):

lol then quest. is wrong @DLS until we r given with u = 0 km/h or 0 m/s

OpenStudy (xishem):

You have to make an assumption because you have no other choice. Given the context of the question, it is either impossible to answer, or you have to make an assumption that v_i is 0.

mathslover (mathslover):

right @Xishem

OpenStudy (dls):

okay..

mathslover (mathslover):

sorry ...

OpenStudy (dls):

thanks

mathslover (mathslover):

no problem though

OpenStudy (xishem):

Any other question I've seen similar to this starts off something like "An object initially at rest...". The wording of this question is just bad.

mathslover (mathslover):

right.. initially at rest means u = 0 m/s